OpenStudy (anonymous):

Please help! What is the solution set of the absolute value sentence |3k + 4| = 8? A. {4/3, -4} B. {x|x > 4 or x < -4/5} C. No Solution D. {x|-11 < x < -7} E. All Real Numbers

OpenStudy (anonymous):

same as the question you asked before. Why don't you try to give the answer? We can check it for you. :D

OpenStudy (anonymous):

do you do online school?

OpenStudy (anonymous):

Yes i do online school, it's my first year. And ok CoCo

OpenStudy (anonymous):

I can't make this one out, i have no clue.

OpenStudy (anonymous):

|3k + 4| = 8 3k+4 =8 or 3k+4=-8 for 3k+4=8, 3k+4-4=8-4 3k=4 k=4/3 for 3k+4=-8, 3k+4-4=-8-4 3k=-12 k=-4 I believe that you can find out the answer now :D